OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)? (x − 2)squared + (y + 3)squared = 5 (x + 2)squared + (y − 3)squared = 25 (x − 2)squared + (y + 3)squared = 25 (x − 2)squared − (y + 3)squared = 5

OpenStudy (anonymous):

@Deezzz

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@rvc

OpenStudy (welshfella):

the general form is (x - a)^" + (y - b)^2 = r^2 where (a,b) is the center and r is the radius

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

so you should be able to work out the left side now by plugging in a = 2 and b = -3

OpenStudy (welshfella):

and the radius will be the distance between the center and the point (-2,0)

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw - if you need more help let me know

rvc (rvc):

:) all the best

OpenStudy (anonymous):

@Deezzz

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@triciaal

hero (hero):

In general, the standard form of the equation of a circle is \((x - h)^2 + (y - k)^2 = r^2\). In this case, (h,k) = (2,3) and (x,y) = (-2,0). Plug both points in to the equation, then simplify and solve for r. Next re-write the standard form of the equation, except this time, insert the given center point and the value of r.

OpenStudy (anonymous):

ok ill try, thanks